OpenStudy (anonymous):

to the nearest tenth of a cubic inch,how much volume is added to aspherical balloon whose radius grows from 10 to 20 inches long

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 } \pi \times 20^{3} - \frac{ 4 }{ 3 } \pi \times 10^{3}\]

OpenStudy (kainui):

What's the formula for volume of a sphere? By calculus we could've figured it out, so you have that to look forward to! In the meantime though, the formula for volume of a sphere is: \[V=\frac{ 4 }{ 3 }\pi r^3\] So you can see that if you have the radius, you can find the volume of the balloon when its radius is 10 or 20. So now you can think about it a second, maybe draw a picture or imagine the two shapes. One is bigger than the other. If you could somehow use these two bits of information of volume to find the other volume, all other geometry problems like this become a breeze.